what is the measurement of an angle between the hour and the minute hands if the time is 8:20?
Answer by Alan3354(6092)

what is the measurement of an angle between the hour and the minute hands if the time is 8:20?
----------
The minute hand is 120 degs from up.
The hour hand is 240 degs + 1/3 of the way to 9:00 = 240 + 10
250 - 120 = 130 degrees apart